1) 200.0 g of Ca(C2H3O2)2 was found in the chemistry laboratory.  

    Calculate 


a) . the volume at STP if Ca(C2H3O2)2 is converted to gas.     

b) .  the molarity of a solution made by dissolving 200.0 g of Ca(C2H3O2)2 in 200 mls of water


1
Expert's answer
2021-01-25T04:21:12-0500

 Ca(C2H3O2)2 = 40 + (24 +3 +32) * 2

= 40 + 118 = 158g/mol

mole = mass/molar mass

= 200.0 g/158g/mol

= 1.26mol

  1. volume = mole * molar volume

= 1.26mol * 22.4dm3/mol

= 28.22dm3

2. molarity = mole /volume

= 1,26mol/ 0.2 L

= 0.252mol/L